# Ultrafast Solvent Dynamics Drives the Formation of the Hydrated Electron in Photoexcited Water

### v1

@@ Abstract
The photophysics and photochemistry associated with irradiating UV light in liquid water is central to numerous physical, chemical and biological processes. One of the key events involved in this process is the generation of the hydrated electron. Despite long study from both experimental and theoretical fronts, a unified understanding of the underlying mechanisms associated with the generation of the solvated electron have remained elusive. Here, using excited-state molecular dynamics simulations of condensed phase photoexcited liquid water, we unravel the key sequence of chemical events leading to the creation of the hydrated electron on the excited state. The process begins through the excitation of electrons localized mostly on specific topological defects in water’s hydrogen network which is subsequently followed by two main reaction pathways. The first, leads to the creation of a hydrogen atom culminating in non-radiative decay back to the ground-state within 100 femtoseconds. The second involves a proton coupled electron transfer, giving rise to the formation of the hydronium ion, hydroxyl radical and the hydrated excess electron on the excited-state. This process is facilitated by ultrafast coupled rotational and translational motions of water molecules leading to the formation of water mediated ion-radical pairs in the network. These species can survive on the picosecond timescale and ultimately modulate the emission of visible photons. All in all, our findings provide fresh perspectives into the interpretation of several independent time-dependent spectroscopies measured over the last decades, paving the way for new directions on both theoretical and experimental fronts.
@@ 1 Introduction
The hydrated electron has captivated scientific interest since its discovery nearly 60 years ago1. This chemical species plays a fundamental role in diverse areas spanning radiation chemistry2, 3, DNA damage4, 5, and redox reactions6, 7. One of the most widely employed methods to generate hydrated electrons involve photoexcitation across a broad range of photon energies, from UV to $\gamma$-ray irradiation8, 9, 2. This species has been generated and studied in various aqueous environments, including bulk liquids, clusters in vacuum and at interfaces10, 11, 12, 13, 14, 15. Interestingly, the hydrated electron can be created upon photoexcitation to the first electronic excited state of liquid water16, 17. This is particularly intriguing given that the excitation energy involved ($\sim$8.3 eV)18 is well below the ionization potential of water ($\sim$11 eV)18, 19. Perhaps even more surprising is that experiments have reported electron generation at photon energies as low as 6.5 eV20, 21.
@@
Understanding how hydrated electrons are produced upon excitation to the first absorption band in liquid water requires a detailed characterization and understanding of the initial photoexcitation event. Numerous experiments have shown that excitation at different energies using either single22, 23 or two photon16, 24 within the first broad band in liquid water (approximately 7-8.5 eV), open different pathways in the excited state. However, unraveling excited state dynamics after photoexcitation in liquid water remains a significant challenge due to the extremely short timescales involved (less than 1 ps) and the simultaneous formation of multiple reactive species (e.g. electrons, protons and hydroxyl radicals to name a few)25, 2, 26. While experiments can successfully identify these final products, which molecular species in the hydrogen-bond network get excited and the intermediate steps that occur immediately after photoexcitation in realistic condensed phase conditions, remain largely unresolved.
@@
To address these limitations, a variety of theoretical models and computational techniques have been developed27. These efforts aim to complement experimental results and provide insight into the transient processes underlying electron generation and solvation in water. A central question in this regard is whether the excitation is localized on a single water molecule or delocalized over multiple molecules. Although the dominant view supports localization28, 29, 17, to the best of our knowledge, most previous theoretical studies do not capture the coupling between thermal disorder and condensed phase fluctuations, thereby limiting the generality of their conclusions. In addition, due to the quantum-mechanical nature of the hydrated electron, one conceptual and practical challenge is how to best localize its position within the hydrogen-bond network. As noted by Herbert, who has pioneered the theory and simulations of the hydrated electron in the electronic ground state, most studies rely exclusively on visual inspection of the electronic wavefunction or density which can bias the interpretation27. This aspect also applies to studies of underlying nature of the electronic excitation in condensed liquid water. Importantly, the hydrogen bonding environment surrounding the molecules that are excited, is known to significantly influence both the energy and intensity of electronic transitions30, 31, 32.
@@
Photoionization of water is known to create the solvated electron33, 24, 34. The equilibrium structure of the hydrated electron remains a subject of ongoing debate. Early theoretical models proposed either a cavity model, where the electron resides in a vacuum-like pocket surrounded by waters35, 36, 37 or non-cavity models in which no excluded volume is formed and the electron diffuses to several water molecules38, 39. These earlier methods relied on parameterized interactions between water molecules and the electron, which limits their predictive power. As a result, more accurate quantum chemistry approaches have become popular in recent years. Ab-initio methods have been applied to water and halide-water clusters in vacuum40, 41, 42, but how and if these results can be extrapolated to bulk water is not straightforward. Notably, some of these models failed to accurately predict the coordination number of the hydrated electron in the liquid phase27. Ab-initio bulk water simulations have been also performed for the hydrated electron24, 33, 43. While all of these approaches have provided valuable insights into the properties of the hydrated electron at equilibrium in the electronic ground state, they do not capture the initial stages of its formation. This is due to the fact that in most of these studies, the simulations start with a pre-inserted electron in the electronic ground state and therefore do not capture the events on the excited-state.
@@
In this work, we investigate the excited-state dynamics of neat liquid water following excitation to the first electronic excited state using the Restricted Open Kohn Sham (ROKS) approach44, 45 a technique that has been successfully used to investigate dissociative and charge transfer excitations in closed-shell systems46, 47, 48, 49. By performing simulations on an ensemble of thermally sampled configurations from the electronic ground state, we elucidate new insights into the photophysics and photochemistry of water paying special attention to the important role of fluctuations of the topology of the water hydrogen-bond network. Our study begins by analyzing the fundamental properties of the first excited state in liquid water, addressing the long-debated question of localization. The canonical manner of tackling this in quantum-chemistry is through visual inspection of the electronic wavefunction or electronic density. Here, by using a technique deployed in the quantum physics community used to estimate the degree of delocalization of a quantum state, dubbed as the Inverse Participation Ratio (IPR)50, 51, we directly probe in a fully unsupervised manner the partitioning of the electronic spin density among the photoexcited water molecules. In addition, the analysis of the local environment of the excited water molecules on the red-tail range of the spectrum reveal that most of them reside on topological defects in the hydrogen bond network52, 28, 30 where the asymmetry of the hydrogen bonding also plays a critical role.
@@
Our theoretical predictions provide a coherent and unified understanding of the mechanisms underlying the photochemistry of water drawn from several different experimental time-dependent spectroscopic measurements over the last two decades. Specifically, our simulations provide new insights into the two proposed mechanisms invoked to rationalize the excited-state dynamics of water, namely Hydrogen Atom Transfer (HAT) and Proton Coupled Electron Transfer (PCET). For both HAT and PCET, one fundamental challenge has been identifying which transient species form after absorption of a photon on the excited state itself, in contrast to being formed as a consequence of photodeactivation to the ground-state. Our approach overcomes these hurdles predicting that the HAT mechanism exhibits a higher quantum yield, consistent with experimental observations16. Additionally, we observe the formation of a transient hydronium radical, a species previously proposed but not yet experimentally confirmed53, 54. Importantly, our results also show that this mechanism does not lead to the formation of a hydrated electron in the excited state. On the other hand, for the PCET pathway, we observe the formation and early-time evolution of the hydrated electron before the system relaxes to the ground-state. The creation of the hydrated electron in the excited-state is in turn driven by coupled translational and rotational motions that lead to hydronium ion and hydroxyl radical pairs mediated by solvent as recently suggested by ultrafast electron diffraction measurements34. Finally, we demonstrate that the photon emission by the hydrated electron measured by Tauber et. al.55 is strongly modulated by the extent of the electron localization. These results offer new perspectives into tuning the color of fluorescence emission of excited state hydrated electrons in aqueous systems.
@@ 2.1 Initial photo-absorption in liquid water
Understanding the nature of the initial photo-absorption in liquid water is a critical step toward elucidating the mechanisms that lead to hydrated electron formation, especially at excitation energies within the first absorption band. Figure 1 summarizes our main findings regarding the characteristics of this initial excitation process employing 100 different conformations sampled from well converged deep neural-network based ground-state simulations (see Computational Methods section for details). Figure 1A presents the experimental absorption spectrum compared with the probability density function (PDF) of the excitation energy for the $S_{0}\to S_{1}$ transition, determined using the ROKS method for all the conformations (see the Computational Methods section). Our simulations predict an average excitation energy of 8.1 eV, which is in excellent agreement with the experimental value of 8.3 eV18, 17 and also with Time-Dependent Density Functional Theory (TDDFT), see section S1 in the Supplementary Information.
@@
To quantify the degree of the localization of the first electronic excitation, we partitioned the electronic spin density across individual water molecules of the system and computed the Inverse Participation Ratio56 (IPR). This approach provides an agnostic manner of determining the extent to which the excitation is delocalized (see Analysis Protocols section). As shown in the Figure 1B, the majority of configurations involve a single water molecule in the $S_{0}\to S_{1}$ transition, although a non insignificant fraction exhibits a more delocalized excitation spread over up to five water molecules. These events typically involve a chain of connected water molecules forming a water-wire57, 58, 59, similar to what was recently observed in many-body theory electronic structure calculations in ice60. The typical analysis used to characterize the extent of delocalization rely on visual inspection of molecular orbitals, electronic densities, or spin densities represented as isosurfaces, which can be highly sensitive to the chosen isovalue and thus somewhat arbitrary27. Our approach overcomes these limitations providing a quantitative description of the excitation which allows also for fractional contributions of the number of water molecules. In addition, previous studies have argued that the first transition is localized on a single water, however most of them were limited to isolated monomers or small water clusters28, 30, 31, where the system size inherently restricts delocalization. Other study has employed periodic water systems of varying boxes sizes, inferring localization from the weak dependence of excitation energies on the system size29. While these approaches often produce excitation energies and oscillator strengths that align well with experimental results, they generally rely on limited sampling and thus fail to capture the influence of thermal fluctuations on the excitation process. Understanding the degree of electronic delocalization upon excitation to the first band is particularly important, as one and two photon absorption experiments can access different electronic configurations22, 16, 24. These configurations, in turn, may lead to distinct pathways in the excited-state dynamics of liquid water.
@@
Figure 1: Photo-initial absorption of neat liquid water. Panel A: Experimental absorption spectra in solid red line (reproduced from the ref 17) and the calculated $S_{0}\to S_{1}$ excitation using ROKS for 100 different conformations obtained from the electronic ground state. Panel B: Probability Distribution Function (PDF) of water molecules involved in the excitation using the Inverse Participation Ratio of the spin densities for all the conformations. The insets show two examples involving 1 and 5 water molecules. Panel C: PDF of the different defects in the Hydrogen Bond (HB) Network of all the water molecules involved in the initial excitation. The insets show the molecular geometries of the normal and each type of defect in the HB network.
@@
Our results clearly show that the water molecules most frequently involved in the lowest electronic transition are those found in hydrogen bond (HB) environments with defects, particularly those with at least one missing acceptor hydrogen bond. These include configurations such as AD (a water molecule accepting and donating a single hydrogen bond) and ADD (a water molecule donating two and accepting one hydrogen bond), as illustrated in Figure 1C. This observation is rooted in the nature of the electronic transition, which corresponds to a transition from a non-bonding orbital ($n$) to an anti-bonding orbital ($\sigma^{*}$). In a well-formed HB network, the non-bonding orbitals are stabilized through their participation in hydrogen bonding. This stabilization raises the energy required for excitation28, 30. In contrast, when a water defect does not accept a hydrogen bond, the non-bonding orbitals are no longer stabilized by the field from the nearby proton and therefore remain at higher energy. As a result, the energy gap for the $n\to\sigma^{*}$ transition becomes smaller. These arguments are essentially also at the root of the blue shift of about 0.7 eV observed in the first absorption band of water when comparing the liquid phase to the gas phase17, 18.
@@ 2.2 Photo-chemical mechanism of water deactivation
The excited state dynamics of liquid water reveal two distinct decay pathways: an ultrafast process with an average lifetime of $\sim$25 fs, and a slower one of $\sim$400 fs (see Figure S4 in the Supplementary Information). Our protocol, described in detail in the Computational Methods section, defines the decay time as the instant when the energy gap between the $S_{1}$ and $S_{0}$ states drops below 0.2 eV. This approach has been successfully applied to describe excited state relaxation pathways in other systems61, 62 and, as we will show later, leads to insights that reinforce and enrich the interpretations of existing experimental observations. Before analyzing the full ensemble, we illustrate these mechanisms using representative trajectories in order to build our intuition on the underlying processes.
Figure 2 depicts the two decay pathways: Hydrogen Atom Transfer (HAT, upper panels) and Proton Coupled Electron Transfer (PCET, lower panels). The HAT mechanism exhibits an ultrafast decay. As shown in Figure 2A, which presents the evolution of the $S_{1}$ and $S_{0}$ potential energy surfaces, the system reaches the non-radiative $S_{1}\to S_{0}$ crossing at approximately 10 fs in this trajectory. During the initial few femtoseconds, one water molecule undergoes O–H bond dissociation. This is evident in Figure 2B, which shows the coordination number of the dissociating hydrogen over time which goes from 1 to 0 (see Analysis Protocols section). This behavior is consistent with the strongly dissociative character of the $S_{1}(n\to\sigma^{*})$ potential energy surface and agrees with previous theoretical and experimental findings18, 17, 63. To determine whether the dissociating species is a hydrogen atom ($\mathrm{H}^{\bullet}$) or a proton ($\mathrm{H}^{+}$), we determined the center of the electron and its distance to the dissociating hydrogen (see Analysis Protocols section for details). As shown in Figure 2C, this distance rapidly drops from 1.2 to below 0.2 Å, clearly indicating that the observed process involves hydrogen atom dissociation.
The PCET mechanism displays a slower decay, on the order of hundreds of femtoseconds (see Figure 2D). In fact, 2 out of 100 trajectories remain in the $S_{1}$ excited state for the entire 1 ps simulation. Following photoexcitation of a water molecule, in addition to the OH dissociation, an electron is released into the surrounding hydrogen-bond network. Figure 2E shows the evolution of the gyration radius of the electron over time (see Analysis Protocols section for details), which serves as a proxy for its spatial delocalization. Initially, the electron is highly delocalized, but around 100 fs becomes increasingly localized. This behavior is consistent with experimental observations33, 64, which report a transition from a very delocalized electron (gyration radius above 20 Å) to a more localized wet electron, and finally to the fully hydrated electron in the ground state (gyration radius 2.4 Å).
@@
Using the features described in Figure 2, specifically the gyration radius, distance between the electron center and the closest hydrogen, we can clearly distinguish between the two decay mechanisms. The detailed protocol for this classification is provided in the Analysis Protocols section. While a comprehensive discussion of each mechanism across the full ensemble of trajectories is presented in the following two sections, it is worth noting that our results indicate that approximately 53% of trajectories follow the HAT pathway, while the remaining exhibit the PCET mechanism. Experimental studies report a 45% occurrence of the HAT pathway at 6.7 eV excitation and 70% at 8.4 eV65, 66. Our result lie between these two values, in excellent agreement with the fact that our simulations probe the red edge of the first absorption band (8.1 eV).

### v2

@@ Abstract
The photophysics and photochemistry associated with irradiating UV light in liquid water is central to numerous physical, chemical and biological processes. One of the key events involved in this process is the generation of the hydrated electron. Despite long study from both experimental and theoretical fronts, a unified understanding of the underlying mechanisms associated with the generation of the solvated electron have remained elusive. Here, using excited-state molecular dynamics simulations of condensed phase photoexcited liquid water, we unravel the key sequence of chemical events leading to the creation of the hydrated electron on the excited state. The process begins through the excitation localized mostly on specific topological defects in the hydrogen-bond network of water which is subsequently followed by two main reaction pathways. The first, leads to the creation of a hydrogen atom culminating in non-radiative decay back to the ground-state within 100 femtoseconds. The second involves a proton coupled electron transfer, giving rise to the formation of the hydronium ion, hydroxyl radical and the hydrated excess electron on the excited-state. This process is facilitated by ultrafast coupled rotational and translational motions of water molecules leading to the formation of water mediated ion-radical pairs in the network. These species can survive on the picosecond timescale and ultimately modulate the emission of visible photons. All in all, our findings provide fresh perspectives into the interpretation of several independent time-dependent spectroscopies measured over the last decades, paving the way for new directions on both theoretical and experimental fronts.
@@ 1 Introduction
The hydrated electron has captivated scientific interest since its discovery nearly 60 years ago1. This chemical species plays a fundamental role in diverse areas spanning radiation chemistry2, 3, DNA damage4, 5, and redox reactions6, 7. One of the most widely employed methods to generate hydrated electrons involve photoexcitation across a broad range of photon energies, from UV to $\gamma$-ray irradiation8, 9, 2. This species has been generated and studied in various aqueous environments, including bulk liquids, clusters in vacuum and at interfaces10, 11, 12, 13, 14, 15, 16. Interestingly, the hydrated electron can be created upon photoexcitation to the first electronic excited state of liquid water17, 18. This is particularly intriguing given that the excitation energy involved ($\sim$8.3 eV)19 is well below the ionization potential of water ($\sim$11 eV)19, 20. Perhaps even more surprising is that experiments have reported electron generation at photon energies as low as 6.5 eV21, 22.
@@
Understanding how hydrated electrons are produced upon excitation to the first absorption band in liquid water requires a detailed characterization and understanding of the initial photoexcitation event. Numerous experiments have shown that excitation at different energies using either single23, 24 or two photon17, 25 within the first broad band in liquid water (approximately 7-8.5 eV), open different pathways in the excited state. However, unraveling excited state dynamics after photoexcitation in liquid water remains a significant challenge due to the extremely short timescales involved (less than 1 ps) and the simultaneous formation of multiple reactive species (e.g. electrons, protons and hydroxyl radicals to name a few)26, 2, 27. While experiments can successfully identify these final products, which molecular species in the hydrogen-bond network get excited and the intermediate steps that occur immediately after photoexcitation in realistic condensed phase conditions, remain largely unresolved.
@@
To address these limitations, a variety of theoretical models and computational techniques have been developed28. These efforts aim to complement experimental results and provide insight into the transient processes underlying electron generation and solvation in water. A central question in this regard is whether the excitation is localized on a single water molecule or delocalized over multiple molecules. Although the dominant view supports localization29, 30, 18, to the best of our knowledge, most previous theoretical studies do not capture the coupling between thermal disorder and condensed phase fluctuations, thereby limiting the generality of their conclusions. In addition, due to the quantum-mechanical nature of the hydrated electron, one conceptual and practical challenge is how to best localize its position within the hydrogen-bond network. As noted by Herbert, who has pioneered the theory and simulations of the hydrated electron in the electronic ground state, most studies rely exclusively on visual inspection of the electronic wavefunction or density which can bias the interpretation28. This aspect also applies to studies of underlying nature of the electronic excitation in condensed liquid water. Importantly, the hydrogen bonding environment surrounding the molecules that are excited, is known to significantly influence both the energy and intensity of electronic transitions31, 32, 33.
@@
Photoionization of water is known to create the solvated electron34, 25, 35, 36. The equilibrium structure of the hydrated electron remains a subject of ongoing debate. Early theoretical models proposed either a cavity model, where the electron resides in a vacuum-like pocket surrounded by waters37, 38, 39 or non-cavity models in which no excluded volume is formed and the electron diffuses to several water molecules40, 41. These earlier methods relied on parameterized interactions between water molecules and the electron, which limits their predictive power. As a result, more accurate quantum chemistry approaches have become popular in recent years. Ab-initio methods have been applied to water and halide-water clusters in vacuum42, 43, 44, but how and if these results can be extrapolated to bulk water is not straightforward. Notably, some of these models failed to accurately predict the coordination number of the hydrated electron in the liquid phase28. Ab-initio bulk water simulations have been also performed for the hydrated electron25, 34, 45. While all of these approaches have provided valuable insights into the properties of the hydrated electron at equilibrium in the electronic ground state, they do not capture the initial stages of its formation. This is due to the fact that in most of these studies, the simulations start with a pre-inserted electron in the electronic ground state and therefore do not capture the events on the excited-state.
@@
In this work, we investigate the excited-state dynamics of neat liquid water following excitation to the first electronic excited state using the Restricted Open Kohn Sham (ROKS) approach46, 47 a technique that has been successfully used to investigate dissociative and charge transfer excitations in closed-shell systems48, 49, 50, 51. By performing simulations on an ensemble of thermally sampled configurations from the electronic ground state, we elucidate new insights into the photophysics and photochemistry of water paying special attention to the important role of fluctuations of the topology of the water hydrogen-bond network. Our study begins by analyzing the fundamental properties of the first excited state in liquid water, addressing the long-debated question of localization. The canonical manner of tackling this in quantum-chemistry is through visual inspection of the electronic wavefunction or electronic density. Here, using the Inverse Participation Ratio (IPR)52, 53 which has been previously applied to study the extent of delocalization of exciton wavefunctions in molecular systems54, we directly probe the partitioning of the electronic spin density among the photoexcited water molecules. In addition, the analysis of the local environment of the excited water molecules on the red-tail range of the spectrum reveal that most of them reside on topological defects in the hydrogen bond network55, 29, 31 where the asymmetry of the hydrogen bonding also plays a critical role.
@@
Our theoretical predictions provide a coherent and unified understanding of the mechanisms underlying the photochemistry of water drawn from several different experimental time-dependent spectroscopic measurements over the last two decades. Specifically, our simulations provide new insights into the two proposed mechanisms invoked to rationalize the excited-state dynamics of water, namely Hydrogen Atom Transfer (HAT) and Proton Coupled Electron Transfer (PCET). For both HAT and PCET, one fundamental challenge has been identifying which transient species form after absorption of a photon on the excited state itself, in contrast to being formed as a consequence of photodeactivation to the ground-state. Our approach overcomes these hurdles predicting that the HAT mechanism exhibits a higher quantum yield, consistent with experimental observations17. Additionally, we observe the formation of a transient hydronium radical, a species previously proposed but not yet experimentally confirmed56, 57. Importantly, our results also show that this mechanism does not lead to the formation of a hydrated electron in the excited state. On the other hand, for the PCET pathway, we observe the formation and early-time evolution of the hydrated electron before the system relaxes to the ground-state. The creation of the hydrated electron in the excited-state is in turn driven by coupled translational and rotational motions that lead to hydronium ion and hydroxyl radical pairs mediated by solvent as recently suggested by ultrafast electron diffraction measurements35. Finally, we demonstrate that the photon emission by the hydrated electron measured by Tauber et. al.58 is strongly modulated by the extent of the electron localization. These results offer new perspectives into tuning the color of fluorescence emission of excited state hydrated electrons in aqueous systems.
@@ 2.1 Initial photo-absorption in liquid water
Understanding the nature of the initial photo-absorption in liquid water is a critical step toward elucidating the mechanisms that lead to hydrated electron formation, especially at excitation energies within the first absorption band. Figure 1 summarizes our main findings regarding the characteristics of this initial excitation process employing 100 different conformations sampled from well converged deep neural-network based ground-state simulations (see Computational Methods section for details). Figure 1A presents the experimental absorption spectrum compared with the probability density function (PDF) of the excitation energy for the $S_{0}\to S_{1}$ transition, determined using the ROKS method for all the conformations (see the Computational Methods section). Our simulations predict an average excitation energy of 8.1 eV, which is in excellent agreement with the experimental value of 8.3 eV19, 18 and also with Time-Dependent Density Functional Theory (TDDFT) (see section S1 in the Supplementary Information).
@@
To quantify the degree of the localization of the first electronic excitation, we partitioned the electronic spin density across individual water molecules of the system and computed the Inverse Participation Ratio59 (IPR). This approach provides an agnostic manner of determining the extent to which the excitation is delocalized (see Analysis Protocols section). As shown in the Figure 1B, the majority of configurations involve a single water molecule in the $S_{0}\to S_{1}$ transition, although a non insignificant fraction exhibits a more delocalized excitation spread over up to five water molecules. These events typically involve a chain of connected water molecules forming a water-wire60, 61, 62, similar to what was recently observed in many-body theory electronic structure calculations in ice63. The typical analysis used to characterize the extent of delocalization rely on visual inspection of molecular orbitals, electronic densities, or spin densities represented as isosurfaces, which can be highly sensitive to the chosen isovalue and thus somewhat arbitrary28. Our approach overcomes these limitations providing a quantitative description of the excitation which allows also for fractional contributions of the number of water molecules. In addition, previous studies have argued that the first transition is localized on a single water, however most of them were limited to isolated monomers or small water clusters29, 31, 32, where the system size inherently restricts delocalization. Other studies have employed periodic water systems of varying boxes sizes, inferring localization from the weak dependence of excitation energies on the system size30. While these approaches often produce excitation energies and oscillator strengths that align well with experimental results, they generally rely on limited sampling and thus fail to capture the influence of thermal fluctuations on the excitation process. Understanding the degree of electronic delocalization upon excitation to the first band is particularly important, as one and two photon absorption experiments can access different electronic configurations23, 17, 25. These configurations, in turn, may lead to distinct pathways in the excited-state dynamics of liquid water.
@@
Figure 1: Photo-initial absorption of neat liquid water. Panel A: Experimental absorption spectra in solid red line (reproduced from the ref 18) and the calculated $S_{0}\to S_{1}$ excitation using ROKS for 100 different conformations obtained from the electronic ground state. Panel B: Probability Distribution Function (PDF) of water molecules involved in the excitation using the Inverse Participation Ratio of the spin densities for all the conformations. The insets show two examples involving 1 and 5 water molecules. Panel C: PDF of the different defects in the Hydrogen Bond (HB) Network of all the water molecules involved in the initial excitation. The insets show the molecular geometries of the normal and each type of defect in the HB network.
@@
Our results clearly show that the water molecules most frequently involved in the lowest electronic transition are those found in hydrogen bond (HB) environments with defects, particularly those with at least one missing acceptor hydrogen bond. These include configurations such as AD (a water molecule accepting and donating a single hydrogen bond) and ADD (a water molecule donating two and accepting one hydrogen bond), as illustrated in Figure 1C. This observation is rooted in the nature of the electronic transition, which corresponds to a transition from a non-bonding orbital ($n$) to an anti-bonding orbital ($\sigma^{*}$). In a well-formed HB network, the non-bonding orbitals are stabilized through their participation in hydrogen bonding. This stabilization raises the energy required for excitation29, 31. In contrast, when a water defect does not accept a hydrogen bond, the non-bonding orbitals are no longer stabilized by the field from the nearby proton and therefore remain at higher energy. As a result, the energy gap for the $n\to\sigma^{*}$ transition becomes smaller. These arguments are essentially also at the root of the blue shift of about 0.7 eV observed in the first absorption band of water when comparing the liquid phase to the gas phase18, 19.
@@ 2.2 Photo-chemical mechanism of water deactivation
The excited state dynamics of liquid water reveal two distinct decay pathways: an ultrafast process with an average lifetime of $\sim$25 fs, and a slower one of $\sim$400 fs (see Figure S4 in the Supplementary Information). Our protocol, described in detail in the Computational Methods section, defines the decay time as the instant when the energy gap between the $S_{1}$ and $S_{0}$ states drops below 0.2 eV. This approach has been successfully applied to describe excited state relaxation pathways in other systems64, 65 and, as we will show later, leads to insights that reinforce and enrich the interpretations of existing experimental observations. Before analyzing the full ensemble, we illustrate these mechanisms using representative trajectories in order to build our intuition on the underlying processes.
Figure 2 depicts the two decay pathways: Hydrogen Atom Transfer (HAT, upper panels) and Proton Coupled Electron Transfer (PCET, lower panels). The HAT mechanism exhibits an ultrafast decay. As shown in Figure 2A, which presents the evolution of the $S_{1}$ and $S_{0}$ potential energy surfaces, the system reaches the non-radiative $S_{1}\to S_{0}$ crossing at approximately 10 fs in this trajectory. During the initial few femtoseconds, one water molecule undergoes O–H bond dissociation. This is evident in Figure 2B, which shows the coordination number of the dissociating hydrogen over time which goes from 1 to 0 (see Analysis Protocols section). This behavior is consistent with the strongly dissociative character of the $S_{1}(n\to\sigma^{*})$ potential energy surface and agrees with previous theoretical and experimental findings19, 18, 66. To determine whether the dissociating species is a hydrogen atom ($\mathrm{H}^{\bullet}$) or a proton ($\mathrm{H}^{+}$), we determined the center of the electron and its distance to the dissociating hydrogen (see Analysis Protocols section for details). As shown in Figure 2C, this distance rapidly drops from 1.2 to below 0.2 Å, clearly indicating that the observed process involves hydrogen atom dissociation.
The PCET mechanism displays a slower decay, on the order of hundreds of femtoseconds (see Figure 2D). In fact, 2 out of 100 trajectories remain in the $S_{1}$ excited state for the entire 1 ps simulation. Following photoexcitation of a water molecule, in addition to the OH dissociation, an electron is released into the surrounding hydrogen-bond network. Figure 2E shows the evolution of the gyration radius of the electron over time (see Analysis Protocols section for details), which serves as a proxy for its spatial delocalization. Initially, the electron is highly delocalized, but around 100 fs becomes increasingly localized. This behavior is consistent with experimental observations34, 67, which report a transition from a very delocalized electron (gyration radius above 20 Å) to a more localized wet electron, and finally to the fully hydrated electron in the ground state (gyration radius 2.4 Å).
@@
Using the features described in Figure 2, specifically the gyration radius, distance between the electron center and the closest hydrogen, we can clearly distinguish between the two decay mechanisms. The detailed protocol for this classification is provided in the Analysis Protocols section. While a comprehensive discussion of each mechanism across the full ensemble of trajectories is presented in the following two sections, it is worth noting that our results indicate that approximately 53% of trajectories follow the HAT pathway, while the remaining exhibit the PCET mechanism. Experimental studies report a 45% occurrence of the HAT pathway at 6.7 eV excitation and 70% at 8.4 eV68, 69. Our result lie between these two values, in excellent agreement with the fact that our simulations probe the red edge of the first absorption band (8.1 eV).